# Mini-batch Estimation for Deep Cox Models: Statistical Foundations and Practical Guidance

### v2

@@ 3.2 Impact of Batch Size on Cox-NN Training
The previous section focuses on the statistical properties of the mb-MPLE $\tilde{f}^{(s)}_{n}$, the global minimizer of (3.2). In practice, searching for the mb-MPLE through SGD is challenging since the NN training is a highly non-convex optimization problem and can be affected by many factors (Li et al., 2018). It has been noticed that the ratio of the learning rate to the batch size $\gamma/s$ is a key factor that determines SGD dynamics (Goyal et al., 2017; Jastrzebski et al., 2017). Goyal et al. (2017) demonstrates that keeping $\gamma/s$ constant makes the SGD training process almost unchanged on a broad range of batch sizes, which is known as the linear scaling rule. This rule guides hyperparameter tuning in NN such that we can fix either $\gamma$ or $s$ and only tune the other parameter to optimize the behavior of SGD.
It remains unclear whether the linear scaling rule can be applied to Cox-NN training, as the population loss $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ in Cox-NN depends on the batch size. When applying SGD in training a neural network $f_{\theta}$ with a specified structure, we treat the parameter $\theta$ as finite-dimensional. Existing theoretical work focused on the population loss $\mathbb{E}[L(\theta)]$ which is invariant to different batch sizes and assumed that the Hessian equals the covariance of gradient at the truth, that is $\nabla^{2}_{\theta}\mathbb{E}[L(\theta)]|_{\theta=\theta_{0}}=\mathbb{V}[\nabla_{\theta}L(\theta)]|_{\theta=\theta_{0}}$ (Jastrzebski et al., 2017; Xie et al., 2020). These key properties of the objective function are violated for the population loss $\mathbb{E}[L^{(s)}_{Cox}$] under the Cox model, as shown in the following theorem.
@@
For $i$th individual loss $L_{i}$, the property $\nabla^{2}_{\theta}\mathbb{E}[L_{i}(\theta)]|_{\theta=\theta_{0}}=\mathbb{V}[\nabla_{\theta}L_{i}(\theta)]|_{\theta=\theta_{0}}$ immediately implies $\nabla^{2}_{\theta}\mathbb{E}[\frac{1}{s}\sum_{i=1}^{s}L_{i}(\theta)]|_{\theta=\theta_{0}}=s\mathbb{V}[\nabla_{\theta}\frac{1}{s}\sum_{i=1}^{s}L_{i}(\theta)]|_{\theta=\theta_{0}}$ for any mini-batch of $s$ i.i.d. samples. The equality (3.4) indicates this batch-level relation still holds for the negative-log-partial likelihood, which is useful to study the SGD dynamic (Jastrzebski et al., 2017; Xie et al., 2020). Moreover, the inequality (3.5) and Remark 1 together depict the properties of $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$. The global minimizer of $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ is always $\theta_{0}$, which does not depend on the batch size $s$, while the local convexity of $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ increases when $s$ doubles. Figure 1(a) presents an illustrative picture of $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ showing these two properties.
@@ 4.1 Offline Cox Regression
where $I(\theta_{0})$ is the information matrix of $\theta_{0}$. This implies that $H_{s}^{-1}$ decreases towards its lower bound $I(\theta_{0})^{-1}$ as $s$ continues to double, suggesting that the mb-MPLE is less efficient than the MPLE, whose asymptotic variance is $I(\theta_{0})^{-1}$. On the other hand, when the batch size is large and $H_{s}^{-1}$ is approaching $I(\theta_{0})^{-1}$, the efficiency gain from doubling the batch size would diminish, which has been empirically reported in Tarkhan & Simon (2024).
@@ 4.2 Online Cox Regression
The strong convexity of the objective function has been widely used to establish the fast convergence of SGD algorithm to the global minimizer (Ruppert, 1988; Polyak & Juditsky, 1992; Moulines & Bach, 2011; Toulis & Airoldi, 2017). It requires the function to grow as fast as a quadratic function. Specifically, a function $h(x)$ is strongly convex over a domain $\mathcal{X}$ if there exists a constant $\mu>0$ such that its Hessian satisfies $\nu^{T}\nabla^{2}_{x}h(x)\nu\geq\mu>0$ for any $x\in\mathcal{X}$ and vector $\nu\in\{\nu:\lVert\nu\rVert_{2}=1\}$, where $\lVert\cdot\rVert_{2}$ denotes the Euclidean norm. In our case, the population loss of online Cox regression $\mathbb{E}[L^{(s)}_{Cox}(\theta)]=\mathbb{E}\left[-\frac{1}{s}\sum_{i:D_{i}\in D(s)}\Delta_{i}\log\frac{\exp(\theta^{T}X_{i})}{\sum_{j:D_{j}\in D(s)}I(T_{j}\geq T_{i})\exp(\theta^{T}X_{j})}\right]$ is not strongly convex globally over $\mathbb{R}^{p}$. For example, when $p=1$, it is straightforward to verify that $\mathbb{E}[L^{(s)}_{Cox}(\theta)]=O(\theta)$ and the Hessian vanishes when $\theta\to\infty$, hence the global strong convexity does not hold. However, we show that $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ is strongly convex within any ball centered at the origin that contains the true parameter $\theta_{0}$. This local strong convexity facilitates the application of existing SGD convergence results to the Cox regression setting.
@@
where $\Pi_{\mathbb{R}^{p}_{B}}$ is the orthogonal projection operator on the ball $\mathbb{R}^{p}_{B}:=\{\theta\in\mathbb{R}^{p}:\lVert\theta\rVert_{2}\leq B\}$ (Moulines & Bach, 2011). The projection step keeps iterates within the area where the local strong convexity of $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ holds. By applying Theorem 2 of Moulines & Bach (2011), we obtain the following non-asymptotic result with respect to the iteration step $t$.
@@
The high-probability error bound for the projected SGD estimator is also derived and presented in the Appendix. Additionally, we show that the running average SGD (ASGD) estimator $\bar{\hat{\theta}}_{t}=\frac{1}{t}\sum_{k=1}^{t}\hat{\theta}_{k}$ (no projection), achieves the optimal $O(t^{-1})$ rate and $\sqrt{t}(\bar{\hat{\theta}}_{t}-\theta_{0})$ is asymptotically normal for $\gamma_{t}=\frac{C}{t^{\alpha}}$ with $\alpha\in(0.5,1)$ and $C>0$. This is useful for uncertainty quantification in online Cox regression (see Appendix).
@@ 5.3 Linear Scaling Rule for Cox Neural Network
where 8.6-8.6 is used to center $\mathbb{E}[f_{0}(X)]$ towards zero. A Cox-NN is fitted under different SGD learning rates and batch sizes. The full negative log-partial likelihood $L_{Cox}^{(N_{test})}$ was calculated on the same test data to reflect the training history. Identical to the spirit of Goyal et al. (2017), our goal is to match the test errors across batch sizes by only adjusting the learning rate. Figure 3 shows that the linear scaling rule still holds in Cox-NN, especially when batch size is large (the three curves of batch size s=128, 256, 512s=128,\ 256,\ 512 are overlapped with each other). The differences between the curves (e.g., batch sizes 32 and 64) are due to the convexity change of the loss function when performing SGD with different batch sizes.
@@ 7 Discussion
The partial likelihood in the Cox model essentially models the rank of event times through a Plackett-Luce (PL) model (Plackett, 1975; Luce, 1959) with censored outcome data. PL model with neural networks has been widely used in various tasks, such as learning-to-rank (LTR) (Cao et al., 2007) and contrastive learning (Chen et al., 2020). These applications are equivalent to applying Cox-NN to time-to-event data without right-censoring. Thus, our results and discussions can be extended to these tasks as a potential direction for future research.
Data Availability
The AREDS data used in this work are available from the online repository dbGaP (accession: phs000001.v3.p1).

### v3

@@ 3.2 Impact of Batch Size on Cox-NN Training
The previous section focuses on the statistical properties of the mb-MPLE $\tilde{f}^{(s)}_{n}$, the global minimizer of (3.2). In practice, searching for the mb-MPLE through SGD is challenging since the NN training is a highly non-convex optimization problem and can be affected by many factors (Li et al., 2018). It has been noticed that the ratio of the learning rate to the batch size $\gamma/s$ is a key factor that determines SGD dynamics (Goyal and others, 2017; Jastrzebski et al., 2017). Goyal and others (2017) demonstrates that keeping $\gamma/s$ constant makes the SGD training process almost unchanged on a broad range of batch sizes, which is known as the linear scaling rule. This rule guides hyperparameter tuning in NN such that we can fix either $\gamma$ or $s$ and only tune the other parameter to optimize the behavior of SGD.
It remains unclear whether the linear scaling rule can be applied to Cox-NN training, as the population loss $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ in Cox-NN depends on the batch size. When applying SGD in training a neural network $f_{\theta}$ with a specified structure, we treat the parameter $\theta$ as finite-dimensional. Existing theoretical work focused on the population loss $\mathbb{E}[L(\theta)]$ which is invariant to different batch sizes and assumed that the Hessian equals the covariance of gradient at the truth, that is $\nabla^{2}_{\theta}\mathbb{E}[L(\theta)]|_{\theta=\theta_{0}}=\mathbb{V}[\nabla_{\theta}L(\theta)]|_{\theta=\theta_{0}}$ (Jastrzebski et al., 2017; Xie and others, 2020). These key properties of the objective function are violated for the population loss $\mathbb{E}[L^{(s)}_{Cox}$] under the Cox model, as shown in the following theorem.
@@
For $i$th individual loss $L_{i}$, the property $\nabla^{2}_{\theta}\mathbb{E}[L_{i}(\theta)]|_{\theta=\theta_{0}}=\mathbb{V}[\nabla_{\theta}L_{i}(\theta)]|_{\theta=\theta_{0}}$ immediately implies $\nabla^{2}_{\theta}\mathbb{E}[\frac{1}{s}\sum_{i=1}^{s}L_{i}(\theta)]|_{\theta=\theta_{0}}=s\mathbb{V}[\nabla_{\theta}\frac{1}{s}\sum_{i=1}^{s}L_{i}(\theta)]|_{\theta=\theta_{0}}$ for any mini-batch of $s$ i.i.d. samples. The equality (3.4) indicates this batch-level relation still holds for the negative-log-partial likelihood, which is useful to study the SGD dynamic (Jastrzebski et al., 2017; Xie and others, 2020). Moreover, the inequality (3.5) and Remark 1 together depict the properties of $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$. The global minimizer of $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ is always $\theta_{0}$, which does not depend on the batch size $s$, while the local convexity of $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ increases when $s$ doubles. Figure 1(a) presents an illustrative picture of $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ showing these two properties.
@@ 4.1 Offline Cox Regression
where $I(\theta_{0})$ is the information matrix of $\theta_{0}$. This implies that $H_{s}^{-1}$ decreases towards its lower bound $I(\theta_{0})^{-1}$ as $s$ continues to double, suggesting that the mb-MPLE is less efficient than the MPLE, whose asymptotic variance is $I(\theta_{0})^{-1}$. On the other hand, when the batch size is large and $H_{s}^{-1}$ is approaching $I(\theta_{0})^{-1}$, the efficiency gain from doubling the batch size would diminish, which has been empirically reported in Tarkhan and Simon (2024).
@@ 4.2 Online Cox Regression
The strong convexity of the objective function has been widely used to establish the fast convergence of SGD algorithm to the global minimizer (Ruppert, 1988; Polyak and Juditsky, 1992; Moulines and Bach, 2011; Toulis and Airoldi, 2017). It requires the function to grow as fast as a quadratic function. Specifically, a function $h(x)$ is strongly convex over a domain $\mathcal{X}$ if there exists a constant $\mu>0$ such that its Hessian satisfies $\nu^{T}\nabla^{2}_{x}h(x)\nu\geq\mu>0$ for any $x\in\mathcal{X}$ and vector $\nu\in\{\nu:\lVert\nu\rVert_{2}=1\}$, where $\lVert\cdot\rVert_{2}$ denotes the Euclidean norm. In our case, the population loss of online Cox regression $\mathbb{E}[L^{(s)}_{Cox}(\theta)]=\mathbb{E}\left[-\frac{1}{s}\sum_{i:D_{i}\in D(s)}\Delta_{i}\log\frac{\exp(\theta^{T}X_{i})}{\sum_{j:D_{j}\in D(s)}I(T_{j}\geq T_{i})\exp(\theta^{T}X_{j})}\right]$ is not strongly convex globally over $\mathbb{R}^{p}$. For example, when $p=1$, it is straightforward to verify that $\mathbb{E}[L^{(s)}_{Cox}(\theta)]=O(\theta)$ and the Hessian vanishes when $\theta\to\infty$, hence the global strong convexity does not hold. However, we show that $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ is strongly convex within any ball centered at the origin that contains the true parameter $\theta_{0}$. This local strong convexity facilitates the application of existing SGD convergence results to the Cox regression setting.
@@
where $\Pi_{\mathbb{R}^{p}_{B}}$ is the orthogonal projection operator on the ball $\mathbb{R}^{p}_{B}:=\{\theta\in\mathbb{R}^{p}:\lVert\theta\rVert_{2}\leq B\}$ (Moulines and Bach, 2011). The projection step keeps iterates within the area where the local strong convexity of $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ holds. By applying Theorem 2 of Moulines and Bach (2011), we obtain the following non-asymptotic result with respect to the iteration step $t$.
@@
The high-probability error bound for the projected SGD estimator is also derived and presented in the online Supplementary Materials. Additionally, we show that the running average SGD (ASGD) estimator $\bar{\hat{\theta}}_{t}=\frac{1}{t}\sum_{k=1}^{t}\hat{\theta}_{k}$ (no projection), achieves the optimal $O(t^{-1})$ rate and $\sqrt{t}(\bar{\hat{\theta}}_{t}-\theta_{0})$ is asymptotically normal for $\gamma_{t}=\frac{C}{t^{\alpha}}$ with $\alpha\in(0.5,1)$ and $C>0$. This is useful for uncertainty quantification in online Cox regression (see Supplementary Materials).
@@ 5.3 Linear Scaling Rule for Cox Neural Network
where 8.6-8.6 is used to center $\mathbb{E}[f_{0}(X)]$ towards zero. A Cox-NN is fitted under different SGD learning rates and batch sizes. The full negative log-partial likelihood $L_{Cox}^{(N_{test})}$ was calculated on the same test data to reflect the training history. Identical to the spirit of Goyal and others (2017), our goal is to match the test errors across batch sizes by only adjusting the learning rate. Figure 3 shows that the linear scaling rule still holds in Cox-NN, especially when batch size is large (the three curves of batch size s=128, 256, 512s=128,\ 256,\ 512 are overlapped with each other). The differences between the curves (e.g., batch sizes 32 and 64) are due to the convexity change of the loss function when performing SGD with different batch sizes.
@@ 7 Discussion
The partial likelihood in the Cox model essentially models the rank of event times through a Plackett-Luce (PL) model (Plackett, 1975; Luce, 1959) with censored outcome data. PL model with neural networks has been widely used in various tasks, such as learning-to-rank (LTR) (Cao and others, 2007) and contrastive learning (Chen and others, 2020). These applications are equivalent to applying Cox-NN to time-to-event data without right-censoring. Thus, our results and discussions can be extended to these tasks as a potential direction for future research.